OpenStudy (anonymous):

There are about 217,000 registered voters in precint A. This is 3 4/9 the number of voters registered in percint B. How many voters are registered in percint B?

OpenStudy (anonymous):

217,000 * 3.44=746,480 in B

OpenStudy (gw2011):

B=63,000 because A is a little over 3 times the registered voters in B. Here is how I calculated it: (3 4/9)(B)=217,000 (31/9)(B)=217,000 B=(217,000)(9/31) B=63,000 registered voters